OpenStudy (anonymous):

The perimeter of a rectangle is 162 ft. The ratio of the length to the width is 5:4. Find the length of the rectangle.

OpenStudy (anonymous):

45ft.

OpenStudy (anonymous):

first find the constant 2(5x+4x)=162 then *5

OpenStudy (johnweldon1993):

2L + 2W = 162 4L = 5W L = \(\large \frac{5}{4}W\) Substitute that into the 1st equation we have \[\large 2(\frac{5}{4}W) + 2W = 162\] \[\large \frac{10}{4}W + 2W = 162\] \[\large \frac{10 + 8}{4}W = 162\] \[\large 18W = 648\] \[\large W = 36\] If W = 36...then 2L + 2(36) = 162 2L + 72 = 162 2L = 90 L = 45 Now...is \[\large \frac{45}{36} = \large \frac{5}{4}\] Looks like both 45 and 36 are divisible by 9 so...45/9 = 5 and 36/9 = 4 so yes it is correct Length = 45